# MERLIN: Multi-stagE query performance prediction for dynamic paRallel oLap pIpeliNe

### v1

@@ 2.2. Learned Methods
Before the raising of deep learning, some studies used machine learning to predict query performance. Ganapathi et al. take plan-level information into consideration and use KCCA (Bach and Jordan, 2003) to predict multiple performance metrics for a single query(Ganapathi et al., 2009). This method can support both short and long–time-running queries (Huang et al., 2023). Li et al. (Li et al., 2012a) train a boosted regression tree for every operator and estimate the cost of the whole plan by scaling and adding the operators’ cost. Akdere et al. proposed two QPP models (Akdere et al., 2012a). The first is a plan-level model which takes the input of the query plan information including numbers and cardinality of each operator and predicts the overall query performance. The second is an operator-level model which predicts the operator’s cost based on the feature of the operator itself and the cost of its children operators. Then they combined these two methods by using the plan-level model to deal with the materialization of the sub-query plan and the operator-level model to others, in order to avoid the high error of the operator-level models on the materialization sub-plan.
@@
In the area of deep learning-based QPP methods, Marcus et al. proposed NEO (Marcus et al., 2019) and QPPNet (Marcus and Papaemmanouil, 2019). The NEO (Marcus et al., 2019) captures relationships between physical operators and their children operators, they do not effectively capture resource competition among operators across different sub-trees. The QPPNet employs one deep learning-based latency model for each operator. Each model accepts the latency, data, and query features of its children operators and other features and then predicts the latency of its operator. Then it dynamically combines the networks corresponding to all operators based on the structure of the query plan tree and passes data features, query features, and latency information between operators to estimate the overall query latency. It’s a flexible method that can adapt to various tree-structured query plans. However, it has a very high cost to build the network for each plan and an insufficient training throughput due to the heterogeneous network architecture makes it hard to take full advantage of GPU’s data parallel capabilities.
@@
Ji Sun et al. proposed an end-to-end learned cost estimator (Sun and Li, 2019) which includes carefully designed complete query representation and feature engineering by vectorizing encoded physical query plan trees containing predicate information and uses a tree-LSTM model to perform end-to-end prediction of query costs. This method is also designed to tree-shaped physical plan.
@@
Yue Zhao et al. proposed QueryFormer (Zhao et al., 2022), a variant of Transformer that can capture tree-structure data modal more effectively, which can be used for tasks like QPP.
@@
Lin Ma et al. proposed MB2 (Ma et al., 2021), a two-level QPP method that considers concurrent operators. They first build a lightweight resource cost model for each operator, then use an interference model that takes inputs of the estimated cost vector of operators and predicts the query performance. Compared to previous methods, this method is more suitable for OLAP databases that involve parallel query plans. However, the interference model didn’t use any special design to capture the resource competition between operators.
@@
Xuanhe et al. proposed a QPP method for concurrent queries using graph embedding (Zhou et al., 2020). They added edges for operators that have resource sharing or competition between query plans of concurrent queries to construct a DAG. Based on resource competition and data sharing, numerical features are added to the edges. Then a graph embedding model is employed to capture the information among all queries and predict the query performance.
@@
Benjamin et al. proposed a zero-shot cost (latency) prediction method (Hilprecht and Binnig, 2022) to achieve robustness on various data, workloads, and physical designs of databases by carefully designing query representation and using GNN to perform prediction.
@@
Li Yan et al. proposed RAAL (Li et al., 2022), a resource-aware deep cost model, which proposed a node-level resource-aware attention model to predict query latency.
@@
However, none of these studies considered the resource competition within a single parallel query plan and the dynamic modification of pipelines in newborn OLAP databases such as ClickHouse.
3. Overview
@@
Figure 1. Overview of MERLIN, which shows the process of data collection, training, and prediction.
@@
As introduced in section 1, MERLIN decouples the cost prediction process of operators and the whole query. Besides the OCPs, such workflow requires the method to collect detailed runtime resources-cost, time-cost information, and operators’ features to train those OCPs. It also includes a multi-stage model to summarize each operator’s cost of the query plan and predict the total query latency. As shown in Figure 1, MERLIN contains three stages:
  • Stage.1
@@
    Operator Cost Prediction. As the Stage 1, it responds to estimate each SIMD operator’s cost by the OPCs, a set of lightweight neural networks responsible for predicting the corresponding operators’ resource costs. Each model is trained on the operator cost data collected by the Runtime Tracker. They take the system overhead and operators’ calling information as input and output the predicted cost vector of operators. Note that the cost vector is not limited to latency only but includes the hardware resource overhead the operator is expected to consume without considering the impact of other concurrent operators. For a given database software and hardware environment, compared to end-to-end models, OCPs can provide better generalization performance across different query plans. Finally, the cost vectors and vertice features are mapped by FFN separately and added together for later cost calibration.
  • Stage.2
@@
    Cost Calibration. In Stage 2, the Cost Calibrator first uses the Graph Constructor to merge execution paths collected in the probe execution mode of the Runtime Tracker into a data-flow tree as a rich expression of the query plan. Then it uses an attention mechanism to adjust each operator’s resource competition weight based on both the meta-competition matrix and the data-flow tree. After that, the Cost Calibrator uses a GAT to calibrate each operator’s cost vector based on the DAG and predicted operators’ costs during parallel pipeline execution.
  • Stage.3
@@
    Query Performance Prediction. In Stage 3, we combine the calibrated cost vectors and the structure of the data-flow tree into a vector tree and use a TCN model to summarize every operator’s cost vector from the bottom up and predicate the total latency of the whole query. It should be noted that the entire process during Stages 2 and 3 is differentiable, thus we can pass the gradient back to the GAT network and train the GAT and TCN models together.
@@
To support the three stages above, MERLIN also includes a Runtime Tracker, which can collect necessary data for training and inference. The Runtime Tracker made some modifications to the common components like the physical operators, data chunks, and executors. We will introduce it in detail in section 4.
@@
In the rest of this section, we will introduce MERLIN’s Runtime Tracker, and the three stages will be introduced in section 5.
4. Runtime Tracker
4.1. Motivation
To develop a physical plan-level performance predictor for high-performance OLAP DBMSs, we need to collect detailed information to train and evaluate the predictor. However, such DBMS’s parallel and dynamic execution mechanisms prevent us from using tools like ‘EXPLAIN’ to support query performance prediction for the following reasons.
Figure 2. Illustration of dynamic modification. The Join operator with a red background is modified from hash join to merge join and the MergeSort operators are added to the DAG in the probe execution phase.
To make an accurate query performance prediction, it is necessary to acquire detailed operators’ runtime information, such as their resource cost and runtime calling parameters. It would be hard to match each operator’s information with the pipeline structure given by the ‘EXPLAIN’ SQL command since most of them employ dynamic execution mechanisms that allow inserting and modifying operators including adding new execution paths and changing the algorithm used by operators such as sort and join. Taking ClickHouse as an example, it involves dynamic modifications of the pipeline structure and dynamically decides on physical operators’ calling parameters, such as the join algorithm parameter of the JoiningTransform operator, which is hard to foresee during performance prediction.
For the first type of action, assuming the DBMS finds that it mistakenly selected merge join instead of hash join algorithm, it will actively modify the join algorithm. Some DBMSs such as ClickHouse achieve this by passing different types of algorithm parameters to the Join operator, the instance of the operator remains the same, but the operator will have different features and costs for most of the execution, which cannot be obtained by simply using the static pipeline. Other DBMSs such as SparkSQL change the algorithm by replacing the corresponding physical operator. In both situations, relying solely on the information provided in a static pipeline is unreliable.
For the second type of action, taking ClickHouse as an example, assuming that it decides to switch from internal sorting to external sorting, it will involve adding and replacing operators in the pipeline. Other DBMSs like SparkSQL with the AQE mechanism can even adjust the join order during execution. In such a situation, the execution paths of chunks can be completely changed.
Therefore, simply extracting static information from the planner is not enough, the runtime information including calling parameters, pipeline modifications, and chunks execution information is necessary for precise performance prediction. However, the pipeline exported by the ‘EXPLAIN’ statement not only does not contain any of those modification details, but it also cannot provide any runtime information of operators and chunks. Thus, it is difficult to use for QPP. Although it is impossible to obtain all pipeline modifications during the whole execution process while predicting query performance, inspired by ClickHouse’s probe phase (Schulze et al., [n.d.]), we developed a probe execution mode to launch, track, and collect the first several chunks’ execution information rapidly and obtain the pipeline structure, operators’ runtime features and pipeline modifications happened on the starting of the execution.
Considering the dynamic pipeline execution mechanism, those chunks can be generated by different execution paths that correspond to different features and costs. Thus, it would be a time-space graph with unaligned time steps. To collect that information, MERLIN either embeds a variable number of chunks’ features from unaligned time slices into each operator of the pipeline or separates those chunks’ execution paths by constructing and processing a data flow tree. The former solution needs time-space graph convolution on a graph with many padded zero features, which could lead to resource waste and lower accuracy. Therefore, we chose the latter solution, which expands the execution path of multiple chunks involved in the probe execution in the spatial dimension and constructs a data-flow tree by combining them as the representation of the dynamic pipeline structure.
Besides, since MERLIN is a multi-stage method that decouples the operator cost prediction, cost calibration, and query performance prediction, it also has to collect the training data of OCPs separately to avoid interference from other parallel operators.
With all those reasons above, we designed the Runtime Tracker to collect detailed runtime information to support more accurate QPP. For the convenience of discussion, we take ClickHouse, a representative high-performance OLAP DBMS, as an example to introduce the Runtime Tracker’s design in the following subsection. Note that such design is portable between DBMSs since high-performance OLAP DBMSs have similar execution mechanisms such as shown in Table 1, and we assume the target DBMS uses a DAG-based pipeline, which is a more general structure compared with the tree-shape pipeline.
4.2. Design of Runtime Tracker and Data Collection Process
Figure 3. Illustration of the structure of the Runtime Tracker. The Serial Executor in dotted lines needs to be modified from the DBMS’s parallel executor if the DBMS does not provide a serial execution mode like SparkSQL does.
As shown in Figure 3, the Runtime Tracker includes three parts: Executor,
  • Executor. The executor part can usually be regarded as a part of the DBMS’s execution engine. However, in order to obtain the OCPs’ training data without interference from the parallel execution of other operators, MERLIN needs a serial executor. Some DBMSs like SparkSQL provide such a serial execution mode but some like ClickHouse do not. For DBMSs such as ClickHouse, it is necessary to create a serial executor from their parallel executor. Such modification is usually simple since most DBMSs with parallel executors provide a knob for the degree of parallelism (DOP).
  • Execution Mode. MERLIN uses a full-collection mode to collect the training data for Stage 2 and Stage 3. Since MERLIN collects the chunk-level information that cannot be fully obtained during inference, it employs a probe execution mode to obtain a part of the dynamic changing pipeline information.
  • Cost Logger. MERLIN uses very detailed low-level information about the physical pipeline as features, including the operator’s cost, calling parameters, system resource utilization, and chunk execution information. The information is finally collected by Cost Logger, which involves some simple warpers and modifications of the DBMS’s physical operators and data chunk class.
Data for OCPs. The features and execution costs of each single operator that is used to train the OCPs should be collected offline. Since the OCPs aim to predict the cost of each operator without interference from the parallel execution of other operators, its training data should be collected separately by a serial executor. To collect enough data with as few queries as possible, we run Runtime Tracker in full-collection mode, which collects every operator’s runtime features and costs on every data chunk. Although this will cause extra data collecting and training costs compared to using the history workload log, the overhead would not be too high because the OCPs are small MLP networks, and they only need a single-time data collection for a deployed DBMS.
Taking ClickHouse’s code as an example for the convenience of discussion, we wrap the work() function with our Cost Logger to record the operator’s calling parameters, system resource utilization, and resource cost during execution. The Cost Logger uses ioctl library to obtain the CPU instructions, CPU cycles, cache references, and cache misses. It uses the cpuinfo file to obtain the CPU frequency and uses proc/stat to obtain the CPU utilization. The Cost Logger also uses the proc/meminfo file of Linux to measure the memory cost and uses the /proc/PID/io file to measure the I/O read and write cost. After applying these modifications, the serial executor can be regarded as degenerating into a volcanic model. Then we execute workloads and collect each operator’s features and costs of executing each data chunk. The data collection is efficient because there are dozens to hundreds of data chunks for each operator in one query.
Data for Cost Calibrator. As discussed above, MERLIN solves the dynamic modification of pipeline and calling parameters during cost calibration and query performance prediction by expanding the runtime execution paths in the spatial dimension. This is achieved by tracking the execution path of chunks in the probe execution mode. During offline training data collection, MERLIN will set the executor and the tracker to the probe execution mode and make a quick short execution till the first few result chunks are collected. During this procedure, all intermediate chunks’ execution paths and operators’ calls are traced, and the potential dynamic modifications are also logged. The Graph Constructor will use this information in Stage 2 to construct a data-flow tree.
To track and log this information, we modified the code of the chunk and operators. We use the address of each operator instance as its ID to distinguish them. Once chunks are generated or processed, the corresponding operator will append its ID into the transform_addr list of the chunks. During the processing of chunks, Cost Logger would trace the costs and log them. The address of those cost records will also be added to the chunks’ record_addr list. By analyzing those records, MERLIN can construct the data-flow tree for cost calibration and query performance prediction. The detailed construction algorithm will be introduced in  subsection 5.2.
Taking the TPC-DS workload as an example, the probe execution phase’s overhead is about 155ms. With this extra cost, we can obtain the data flow and partial dynamic added execution paths to support more accurate performance prediction. Since ClickHouse is designed for processing complex OLAP queries that take hours to execute, the overhead is relatively small. Besides, this overhead is eliminable by marking those tuples that had been processed during probe execution and reading the result to skip them during the latter query execution.
This approach appears to have a limitation in its inability to capture all dynamic modifications occurring throughout the entirety of query execution, particularly those stemming from system load variations since they are unpredictable. However, it remains effective in detecting modifications arising from estimation errors, such as cardinality errors. These types of adjustments typically occur during the early stages of query execution, which are adequately addressed by the probe phase.
5. MERLIN’s QPP Model
@@
In this section, we introduce MERLIN’s predictor part in the following order: Operator Cost Prediction, Cost Calibration, and Query Performance Prediction.
5.1. Operator Cost Prediction
@@
In this section, we introduce the Operator Cost Predictor (OCP) in detail. To predict high-performance OLAP DBMSs’ query performance, it is necessary to estimate the resource consumption and the current system resource utilization. Knowing that each operator’s resource consumption is the precondition for predicting the resource competition between operators and accurate latency prediction since those DBMSs are designed to maximize resource utilization efficiency with a high degree of parallelism.
@@
The OCPs are designed to predict the resource consumption of the given operator without the interference of other concurrent operators. Therefore, we can decouple and predict the cost of a single operator and the cost of the whole parallel query pipeline separately. This can improve the prediction accuracy for physical operators with SIMD instructions and make it easier for MERLIN to generalize between different query templates to reduce the accuracy degradation problem caused by workload drift. The inputs of the OCPs consist of two parts: Resource Utilization and Operator Features.
@@
  • Resource Utilization: It contains CPU, memory, and I/O utilization before execution, all OCPs have these three features.
@@
There are two types of features, and we encode them separately. For numerical features, we normalize them during encoding to avoid numerical issues. For categorical features including string-type features and other non-numerical features, we use one-hot encoding to represent them.
@@
5.2. Cost Calibration
Cost Calibration is a key procedure of MERLIN due to high-performance OLAP DBMSs employing a high degree of parallelism design. Taking ClickHouse as an example, it uses horizontal and vertical parallel execution. Horizontal parallel execution includes two parts, the operators implemented by SIMD CPU instructions and the operators created multiple times to deal data in multiple threads, just like the sort operator in Figure 2. The former is deals with MERLIN’s decoupling design, and its costs are predicted by OCPs. The latter involves multi-threading which will be represented in the pipeline structure, and its costs should be calibrated by the GAT model introduced in this section. Vertical parallel execution means ClickHouse keeps traversing the pipeline and parallel executes all operators that can be executed. So every operator in the pipeline can be executed anytime once its input port is ready to send data chunks. This design can significantly accelerate execution, but will also cause widespread competition for CPU, memory, and I/O resources in the whole pipeline.
@@ 5.2. Cost Calibration
To calibrate those operators’ costs predicted by the OCPs, MERLIN needs to know the pipeline structure and the resource competition relation among all operators. As we discussed in section 4, the regular ‘EXPLAIN’ tool is not enough for pipelines with dynamic modification mechanisms, and MERLIN’s cost tracker collects more detailed information including tracked execution paths and other physical plan information with probe execution mode. In stage 2 of cost calibration, MERLIN’s Graph Constructor fully uses the tracked execution paths of chunks and constructs a data-flow tree instead of using the DAG pipeline.
@@
To understand such data-flow tree construction better, taking the Figure 2 as an example, it shows the pipeline and chunks during the probe execution. Taking ClickHouse’s official example in their presentations [^4] footnote 3, assuming that ClickHouse dynamically added and replaced some operators such as MergeSort during the probe execution, then all chunks after this modification would be executed through a different path. Suppose that we simply use the modified DAG pipeline as a representation of the query plan. In that case, all execution information before modification has to be dropped since they cannot match the modified pipeline. Such a solution will cause inaccurate estimation. Besides, such a DAG pipeline is also hard to adapt to existing methods for comparison and hard to support other downstream tasks since most of them are designed based on a tree plan that is used by most DBMSs. Therefore, MERLIN takes full use of the information collected during the probe execution mode by merging each chunk’s execution path into a data-flow tree, in which different pipelines exist at the same time.
@@
As shown in algorithm 1, we locate each operator by a chunk’s execution path that starts at the root operator and ends at itself. We first traverse every execution path and split the path into a series of operator’s id at lines 2-3. Then we traverse each operator on that path from top to bottom at line 4. We first check whether the current operator is the root node of the data-flow tree and set it at lines 5-7. If not, then we try to match it in the children of the current node from line 8. If the current node does not have any children, we create a new node for the current operator and connect it as a child of the current node at lines 9-12. Otherwise, we try to match each child node with the current operator at lines 13-15. If we get a match, then we set the matched child node as the current node and start the next round of the loop to match the next operator at line 16. Otherwise, we also create a node with the current operator and connect it as a child at lines 17-20.
@@
By applying such a construction algorithm, the data-flow tree will branch at the red part of the pipeline in Figure 2 since the modified execution path will not match at line 15 of algorithm 1. Therefore, the data flow during probe execution is expanded from the time dimension to the spatial dimension that can express pipelines before and after modification.
@@
The data-flow tree is used as the backbone of the input graph of the Cost Calibrator, where edges representing resource competition are added. Now we explain how the Graph Constructor generates the CPU, memory, and I/O resource competition. We evaluated each operator’s resource consumption to divide them into different resource-bounded types based on both code implementation and the number of their CPU instructions, memory cost, and I/O times while processing the same amount of data. As the results are shown in  Figure 4, an operator can have multiple types of bounded resources. Based on this, the Graph Constructor can build CPU, memory, and I/O meta-resource competition matrices between every distinct operator type separately. Let the number of operator types be $M$. Those matrices can be denoted as $\mathcal{M}^{c}$, $\mathcal{M}^{m}$, $\mathcal{M}^{io}\in R^{M\times M}$. Note that this is only a rough classification, and the resource competition intensity will be adaptively learned based on pipeline structure with an attention mechanism that will be introduced in the next paragraph. For a given data-flow tree $F$ constructed from a pipeline that has $N$ operators, MERLIN treats it as a DAG and expresses it as an adjacency matrix $\mathcal{M}_{F}\in R^{N\times N}$. The Graph Constructor expands those meta resource competition matrices to $\mathcal{M}_{F}^{c},\mathcal{M}_{F}^{m},\mathcal{M}_{F}^{io}\in R^{N\times N}$ with the same dimension by indexing the operator’s resource competition relation by the operators’ order in $\mathcal{M}_{F}$. After being adjusted by an attention mechanism, the results $\mathcal{M^{\prime}}_{F}^{c}$, $\mathcal{M^{\prime}}_{F}^{m}$, $\mathcal{M^{\prime}}_{F}^{io}$ are added to the $\mathcal{M}_{F}$ to obtain the final expression of the pipeline:
(1) $\mathcal{M}_{pipeline}=\mathcal{W}_{F}\mathcal{M}_{F}+\mathcal{W}_{c}\mathcal{M^{\prime}}_{F}^{c}+\mathcal{W}_{m}\mathcal{M^{\prime}}_{F}^{m}+\mathcal{W}_{io}\mathcal{M^{\prime}}_{F}^{io}$

### v2

@@ II-B Learned Methods
Before the rise of deep learning, some studies used machine learning to predict query performance. Ganapathi et al. take plan-level information into consideration and use KCCA [5] to predict multiple performance metrics for a single query[6]. This method can support both short and long-time-running queries [7]. Li et al. [23] train a boosted regression tree for every operator and estimate the cost of the whole plan by scaling and adding the operators’ cost. Akdere et al. proposed two QPP models [8]. The first is a plan-level model that takes the input of the query plan information, including numbers and cardinality of each operator, and predicts the overall query performance. The second is an operator-level model that predicts the operator’s cost based on the feature of the operator itself and the cost of its child operators. Then they combined these two methods by using the plan-level model to deal with the materialization of the sub-query plan and the operator-level model for others, in order to avoid the high error of the operator-level models on the materialization sub-plan.
@@
In the area of deep learning-based QPP methods, Marcus et al. proposed NEO [4] and QPPNet [11]. The NEO [4] captures relationships between physical operators and their child operators, they do not effectively capture resource competition among operators across different sub-trees. The QPPNet employs one deep learning-based latency model for each operator. Each model accepts the latency, data, and query features of its children operators and other features and then predicts the latency of its operator. Then it dynamically combines the networks corresponding to all operators based on the structure of the query plan tree and passes data features, query features, and latency information between operators to estimate the overall query latency. It’s a flexible method that can adapt to various tree-structured query plans. However, it has a very high cost to build the network for each plan, and an insufficient training throughput due to the heterogeneous network architecture makes it hard to take full advantage of the GPU’s data parallel capabilities.
@@
Ji Sun et al. proposed an end-to-end learned cost estimator [13] which includes carefully designed complete query representation and feature engineering by vectorizing encoded physical query plan trees containing predicate information and uses a tree-LSTM model to perform end-to-end prediction of query costs. This method is also designed for a tree-shaped physical plan.
@@
Yue Zhao et al. proposed QueryFormer [10], a variant of Transformer that can capture tree-shaped data modal more effectively, which can be used for tasks like QPP.
@@
Lin Ma et al. proposed MB2 [9], a two-level QPP method that considers concurrent operators. They first build a lightweight resource cost model for each operator, then use an interference model that takes inputs of the estimated cost vector of operators and predicts the query performance. Compared to previous methods, this method is more suitable for OLAP databases that involve parallel query plans. However, the interference model didn’t use any special design to capture the resource competition between operators.
@@
Xuanhe et al. proposed a QPP method for concurrent queries using graph embedding [12]. They added edges for operators that have resource sharing or competition between query plans of concurrent queries to construct a DAG. Based on resource competition and data sharing, numerical features are added to the edges. Then a graph embedding model is employed to capture the information among all queries and predict the query performance.
@@
Benjamin et al. proposed a zero-shot cost (latency) prediction method [24] to achieve robustness on various data, workloads, and physical designs of databases by carefully designing query representation and using GNN to perform prediction.
@@
Li Yan et al. proposed RAAL [17], a resource-aware deep cost model, which proposed a node-level resource-aware attention model to predict query latency.
@@
However, none of these studies considered the resource competition within a single parallel query plan and the dynamic modification of pipelines in high-performance OLAP databases.
III Overview
@@
Figure 1: Overview of CONCERTO, which shows the process of data collection, training, and prediction.
@@
As introduced in section I, CONCERTO decouples the cost prediction process of operators and the whole query. Besides the OCPs, such a workflow requires a method to collect detailed runtime resource costs, time-cost information, and operators’ features to train those OCPs. It also includes a multi-stage model to summarize each operator’s cost of the query plan and predict the total query latency. As shown in Figure 1, CONCERTO contains three stages:
@@
  • Stage 1: Operator Cost Prediction. As the Stage 1, it responds to estimate each SIMD operator’s cost by the OPCs, a set of lightweight neural networks responsible for predicting the corresponding operators’ resource costs. Each model is trained on the operator cost data collected by the Runtime Tracker. They take the system overhead and operators’ calling information as input and output the predicted cost vector of operators. Note that the cost vector is not limited to latency only, but includes the hardware resource overhead that the operator is expected to consume without considering the impact of other concurrent operators. For a given database software and hardware environment, compared to end-to-end models, OCPs can provide better generalization performance across different query plans. Finally, the cost vectors and vertex features are mapped by FFN separately and added together for later cost calibration.
@@
  • Stage 2: Cost Calibration. In Stage 2, the Cost Calibrator first uses the Graph Constructor to merge execution paths collected in the probe execution mode of the Runtime Tracker into a data-flow tree as a rich expression of the query plan. Then it uses an attention mechanism to adjust each operator’s resource competition weight based on both the meta-competition matrix and the data-flow tree. After that, the Cost Calibrator uses a GAT to calibrate each operator’s cost vector based on the DAG and predicted operators’ costs during parallel pipeline execution.
@@
  • Stage 3: Query Performance Prediction. In Stage 3, we combine the calibrated cost vectors and the structure of the data-flow tree into a vector tree and use a TCN model to summarize every operator’s cost vector from the bottom up and predict the total latency of the whole query. It should be noted that the entire process during Stages 2 and 3 is differentiable, thus, we can pass the gradient back to the GAT network and train the GAT and TCN models together.
@@
To support the three stages above, CONCERTO also includes a Runtime Tracker, which can collect necessary data for training and inference. The Runtime Tracker made some modifications to the common components like the physical operators, data chunks, and executors. We will introduce it in detail in section V.
@@
In the rest of this section, we will introduce CONCERTO’s Runtime Tracker, and the three stages will be introduced in section IV.
IV CONCERTO’s QPP Model
@@
In this section, we introduce CONCERTO’s predictor part in the following order: Operator Cost Prediction, Cost Calibration, and Query Performance Prediction.
IV-A Operator Cost Prediction
@@
In this section, we introduce the Operator Cost Predictor (OCP) in detail. To predict high-performance OLAP DBMSs’ query performance, it is necessary to estimate the resource consumption and the current system resource utilization. Knowing that each operator’s resource consumption is the precondition for predicting the resource competition between operators and accurate latency prediction, since those DBMSs are designed to maximize resource utilization efficiency with a high degree of parallelism.
@@
The OCPs are designed to predict the resource consumption of the given operator without the interference of other concurrent operators. Therefore, we can decouple and predict the cost of a single operator and the cost of the whole parallel query pipeline separately. This can improve the prediction accuracy for physical operators with SIMD instructions and make it easier for CONCERTO to generalize between different query templates to reduce the accuracy degradation problem caused by workload drift. The inputs of the OCPs consist of two parts: Resource Utilization and Operator Features.
@@
  • Resource Utilization: It contains CPU, memory, and I/O utilization before execution; all OCPs have these three features.
@@
There are two types of features, and we encode them separately. For numerical features, we normalize them during encoding to avoid numerical issues. For categorical features, including string-type features and other non-numerical features, we use one-hot encoding to represent them.
@@
IV-B Cost Calibration
Cost calibration constitutes a critical component of CONCERTO, particularly for high-performance OLAP DBMSs that extensively employ parallel execution paradigms. These systems typically incorporate three fundamental parallelism strategies: (1) vectorized operator implementations, (2) intra-operator multithreading, and (3) parallel pipeline execution.
CONCERTO’s decoupled architecture handles SIMD-optimized operators through dedicated Operator Cost Predictors (OCPs). For intra-operator parallelism, the pipeline structure explicitly models multithreaded execution, with cost calibration performed by our proposed Graph Attention Network (GAT) model. The most challenging scenario arises from vertical parallel execution, where the DBMS continuously traverses the pipeline and concurrently executes all data-ready operators. While this design maximizes throughput, it introduces system-wide contention for computational resources (CPU cycles), memory bandwidth, and I/O operations across the entire pipeline.
@@ IV-B Cost Calibration
To calibrate those operators’ costs predicted by the OCPs, CONCERTO needs to know the pipeline structure and the resource competition relation among all operators. As we discussed in section V, the regular ‘EXPLAIN’ tool is not enough for pipelines with dynamic modification mechanisms, and CONCERTO’s cost tracker collects more detailed information, including tracked execution paths and other physical plan information, with probe execution mode. In stage 2 of cost calibration, CONCERTO’s Graph Constructor fully uses the tracked execution paths of chunks and constructs a data-flow tree instead of using the DAG pipeline.
@@
To understand such data-flow tree construction better, taking the Figure 3 as an example, it shows the pipeline and chunks during the probe execution. Assuming that the high-performance OLAP DBMS dynamically added and replaced some operators, such as MergeSort, during the probe execution [^5] footnote 4, then all chunks after this modification would be executed through a different path. Suppose that we simply use the modified DAG pipeline as a representation of the query plan. In that case, all execution information before modification has to be dropped since it cannot match the modified pipeline. Such a solution will cause inaccurate estimation. Besides, such a DAG pipeline is also hard to adapt to existing methods for comparison and hard to support other downstream tasks since most of them are designed based on a tree plan that is used by most DBMSs. Therefore, CONCERTO takes full use of the information collected during the probe execution mode by merging each chunk’s execution path into a data-flow tree, in which different pipelines exist at the same time.
@@
As shown in algorithm 1, we locate each operator by a chunk’s execution path that starts at the root operator and ends at itself. We first traverse every execution path and split the path into a series of operators’ IDs at lines 2-3. Then we traverse each operator on that path from top to bottom at line 4. We first check whether the current operator is the root node of the data-flow tree and set it at lines 5-7. If not, then we try to match it in the children of the current node from line 8. If the current node does not have any children, we create a new node for the current operator and connect it as a child of the current node at lines 9-12. Otherwise, we try to match each child node with the current operator at lines 13-15. If we get a match, then we set the matched child node as the current node and start the next round of the loop to match the next operator at line 16. Otherwise, we also create a node with the current operator and connect it as a child at lines 17-20.
@@
By applying such a construction algorithm, the data-flow tree will branch at the red part of the pipeline in Figure 3 since the modified execution path will not match at line 15 of algorithm 1. Therefore, the data flow during probe execution is expanded from the time dimension to the spatial dimension, which can express pipelines before and after modification.
@@
The data-flow tree is used as the backbone of the input graph of the Cost Calibrator, where edges representing resource competition are added. Now we explain how the Graph Constructor generates the CPU, memory, and I/O resource competition. We evaluated each operator’s resource consumption to divide them into different resource-bounded types based on both code implementation and the number of their CPU instructions, memory cost, and I/O times while processing the same amount of data. As the results are shown in  Figure 2, an operator can have multiple types of bounded resources. Based on this, the Graph Constructor can build CPU, memory, and I/O meta-resource competition matrices between every distinct operator type separately. Let the number of operator types be $M$. Those matrices can be denoted as $\mathcal{M}^{c}$, $\mathcal{M}^{m}$, $\mathcal{M}^{io}\in R^{M\times M}$. Note that this is only a rough classification, and the resource competition intensity will be adaptively learned based on the pipeline structure with an attention mechanism that will be introduced in the next paragraph. For a given data-flow tree $F$ constructed from a pipeline that has $N$ operators, CONCERTO treats it as a DAG and expresses it as an adjacency matrix $\mathcal{M}_{F}\in R^{N\times N}$. The Graph Constructor expands those meta resource competition matrices to $\mathcal{M}_{F}^{c},\mathcal{M}_{F}^{m},\mathcal{M}_{F}^{io}\in R^{N\times N}$ with the same dimension by indexing the operator’s resource competition relation by the operators’ order in $\mathcal{M}_{F}$. After being adjusted by an attention mechanism, the results $\mathcal{M^{\prime}}_{F}^{c}$, $\mathcal{M^{\prime}}_{F}^{m}$, $\mathcal{M^{\prime}}_{F}^{io}$ are added to the $\mathcal{M}_{F}$ to obtain the final expression of the pipeline:
$\mathcal{M}_{pipeline}=\mathcal{W}_{F}\mathcal{M}_{F}+\mathcal{W}_{c}\mathcal{M^{\prime}}_{F}^{c}+\mathcal{W}_{m}\mathcal{M^{\prime}}_{F}^{m}+\mathcal{W}_{io}\mathcal{M^{\prime}}_{F}^{io}$ (1)